If the perimeter of a rectangle is 70 inches. What is the perimeter of a rectangle that has twice the width and twice the height of the first rectangle?
Answer by richwmiller(17219)

p=2l+2w=70
2*2l+2*2w=2*(2l+2w)=2*70=140
check
let's test 15 and 20 as the original sides
2*15+2*20=70
2*30+2*40=60+80=140
ok